If the interior angle of a regular n-gon measures 160 degrees, what is n?
Answer by venugopalramana(3286)

FORMULA FOR INTERIOR ANGLE OF N SIDED POLYGON =(2N-4)*90/N=160
160N=180N-360
180N-160N=20N=360
N=360/20=18 SIDES